OpenStudy (anonymous):

how do you solve 3t^6-48t^2=0

OpenStudy (mertsj):

Factor out 3t^2

OpenStudy (mertsj):

3t^2(t^4-16)

OpenStudy (cwrw238):

factor 3t^2(t^4 - 16) = 0 t = 0 , 2 , -2

OpenStudy (mertsj):

Now you have the difference of two squares. Factor that.

OpenStudy (anonymous):

How do you factor that out? I understand how you get that but there should be six answers

OpenStudy (anonymous):

t = 0 , 2 , -2, 2i,-2i

OpenStudy (mertsj):

\[t^4-16=(t^2-4)(t^2+4)=(t-2)(t+2)(t^2+4)\]

OpenStudy (anonymous):

what happened to the 3t^2?

OpenStudy (anonymous):

x=0 double root

OpenStudy (mertsj):

I left it off because I thought you already recognized that it is a factor.

OpenStudy (anonymous):

so how do you get the six answers by factoring is what im confused about

OpenStudy (mertsj):

\[3t^2(t-2)(t+2)(t^2+4)\]

OpenStudy (mertsj):

That is the completely factored form.

OpenStudy (anonymous):

\[x_1=0,x_2=0,x_3=2,x_4=-2,x_5=2i,x_6=-2i\]

OpenStudy (anonymous):

how do you get the six answers though. I know there should be six because the degree is six

OpenStudy (anonymous):

there are 6 answers already!

OpenStudy (anonymous):

How didyou get them though? I see and understand there are six but i dont know HOW you got them

OpenStudy (anonymous):

\[3t^6-48t^2=0\] \[3t^2(t^4-16)=0\] \[3t^2(t-2)(t+2)(t^2+4)=0\]

OpenStudy (mertsj):

If the product of factors is 0 then one of the factors has to be 0. so set each factor to 0 and solve

OpenStudy (phi):

not sure how you are confused. Do you know how to find the *2* roots of \[ t^2= 0\] t= ±0 two solutions, 0 and 0

OpenStudy (anonymous):

\[3t^2=0\] \[t+2=0\] \[t-2=0\] \[t^2+4=0\]

OpenStudy (anonymous):

if a*b=0 then a=0 or b=0

OpenStudy (anonymous):

Okay, now I understand, thankyou so much